Required information The Foundational 15 [LO2-1, LO2-2, LO2-3, LO2-4] [The following information applies to the questions dis6. If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answ7. Assume that Sweeten Company used cost-plus pricing and a markup percentage of 80% of total manufacturing cost) to establis8. What was Sweeten Companys cost of goods sold for March? (Do not round intermediate calculations.) Cost of goods sold
Answer #1

Plantwide overhead rate = (2500*1.4+1500*2.2+25000)/4000 = 7.95 per MH

6) Unit product cost

Job Q
Direct material 8000
Direct labor 7500
Overhead (1700*7.95) 13515
Total Cost 29015
Unit product cost (29015/30) 967

7) calculate following

Job P Job Q
Total price for the Job 52285*1.8 = 94113 29015*1.8 = 52227
Selling price per unit 94113/20 = 4706 1741

8) Cost of goods sold = 52285+29015 = 81300